OpenStudy (anonymous):

While traveling south on an expressway, a car traveling 60 mph (miles per hour), slows down to 30 mph in 12 minutes due to traffic conditions. Calculate the acceleration.

OpenStudy (rock_mit182):

\[a = \frac{ vf-vi }{ t }\]

OpenStudy (rock_mit182):

that makes sense for you ?

OpenStudy (anonymous):

No whats vf or vi?

OpenStudy (rock_mit182):

vf= final velocity vi = initial velocity

OpenStudy (anonymous):

So its -30mph?

OpenStudy (anonymous):

Divided by 1/5 hr?

OpenStudy (anonymous):

So its -150mph?

OpenStudy (rock_mit182):

i guess you have to convert

OpenStudy (anonymous):

What?

OpenStudy (rock_mit182):

i mean the units to m/s

OpenStudy (rock_mit182):

cause an acceleration has units of\[\frac{ m }{ s ^{2} }\]

OpenStudy (anonymous):

Miles o minutes?

OpenStudy (rock_mit182):

do you have the clue i mean in your text book ?

OpenStudy (rock_mit182):

they sometimes give the answer

OpenStudy (anonymous):

It says that if the speed were given in miles per hour and the time in minutes, you could change the minites to a fraction of an hpur and do the problem. THe anwer would be in miles/hr.

OpenStudy (rock_mit182):

look what I wolud do

OpenStudy (rock_mit182):

yes, sorry my bad. i'm used to work with meters per second

OpenStudy (rock_mit182):

so the only thing you have to convert is minuts so you were right

OpenStudy (rock_mit182):

lol, ;) now solve for a

OpenStudy (anonymous):

Which is -150

OpenStudy (rock_mit182):

:) and the units ?

OpenStudy (anonymous):

-150 mi/hr?

OpenStudy (rock_mit182):

ther is something missing, you're two thirds right

OpenStudy (rock_mit182):

check out the time units

OpenStudy (anonymous):

Thats what i dont get?

OpenStudy (rock_mit182):

remember an accelation has units of: \[a =\frac{ longitud }{ time ^{2} }\]

OpenStudy (rock_mit182):